Ozone decomposes to form oxygen molecules and oxygen atoms, O3(g) ? O2(g) + O(g), in the upper atmosphere. The energy of activation for this reaction is 93.1 kJ/mol. At 600 K, the rate constant for this reaction is 3.37
Answer #1

Arrhenius equation: In(k2/k1) = k/mol = 93100 /mol (-Ea/R) x (1/T2-1/T1) Ea = 93.1 R = 8 . 314 J/mol . K Substituting values